OpenStudy (jb1625):

Simplify. –20 + (–2) • (–6) = A. –132 B. –32 C. –8 D. 132

OpenStudy (yolomcswagginsggg):

c

OpenStudy (jb1625):

Thank you

OpenStudy (yolomcswagginsggg):

Np bruh

OpenStudy (defiance):

Yolo

OpenStudy (defiance):

Try and get them to the answer and not give it directly.

OpenStudy (anonymous):

D is the right answer. Please give a medal.

OpenStudy (yolomcswagginsggg):

Fine here

OpenStudy (anonymous):

D is correct.. times a negative by a negative makes posotive

OpenStudy (jb1625):

How do you know D is the right answer?

OpenStudy (yolomcswagginsggg):

20+(−2)(−6)=−20+12=−8

OpenStudy (jb1625):

And Sorry I already gave a medal

OpenStudy (defiance):

-2 x -6 = 12

OpenStudy (defiance):

-20 + 12 = -8

OpenStudy (defiance):

Btw, you can undo the medal.

OpenStudy (jb1625):

Oh cool

OpenStudy (yolomcswagginsggg):

@JB1625 go to mathpapa and put in ur question ;p

OpenStudy (jb1625):

So is it D. or C.?

OpenStudy (defiance):

Negative times a negative makes a positive.

OpenStudy (defiance):

C

OpenStudy (yolomcswagginsggg):

me and @Defiance say its c

OpenStudy (jb1625):

Thank you

OpenStudy (anonymous):

it is D i just did it on a calculator

OpenStudy (defiance):

I guess he got D working left to right, without PEMDAS. Which is order of operation

OpenStudy (jb1625):

Mabye

OpenStudy (anonymous):

-shoots self in head- i was wrong sorry xD

OpenStudy (anonymous):

totally forgot about pemdas

OpenStudy (yolomcswagginsggg):

hold on lemme use a calculator ;p

OpenStudy (defiance):

You guys are doing it without PEMDAS.

OpenStudy (jb1625):

K Thx everyone

OpenStudy (defiance):

Mhm

OpenStudy (anonymous):

c is correct

OpenStudy (defiance):

You're welcome.

OpenStudy (yolomcswagginsggg):

i got -8

OpenStudy (jb1625):

Cool

OpenStudy (defiance):

Yes.

OpenStudy (yolomcswagginsggg):

Alrighty

OpenStudy (defiance):

Need anything else tag me with @defiance.